# Generalized Bayesian deep reinforcement learning

### v1

@@ A.1 Proof of BvM, Theorem 2
$e^{\operatorname{\mathcal{PS}}_{T}\left(\theta_{T}\right)}T^{d/2}z_{T}=\int g_{T}\longrightarrow\int g_{0}=\pi\left(\theta^{*}\right)\frac{(2\pi)^{d/2}}{\left|\operatorname{\Gamma}_{*}\right|^{1/2}},$ (18)
where $\left|\operatorname{\Gamma}_{*}\right|=\left|\operatorname{det}\operatorname{\Gamma}_{*}\right|$, and hence,
$z_{T}\sim\frac{e^{-\operatorname{\mathcal{PS}}_{T}\left(\theta_{T}\right)}\pi\left(\theta^{*}\right)}{\left|\operatorname{\Gamma}_{*}\right|^{1/2}}\left(\frac{2\pi}{T}\right)^{d/2}\quad\text{as $T\rightarrow\infty$}.$
This proves equation (12) in our Theorem 2.
As $T\rightarrow\infty$, for any sequence $b_{T}\rightarrow b_{0}\in\mathbb{R}$, we have $\int\left|b_{T}g_{T}-b_{0}g_{0}\right|\rightarrow 0$, since (as $\int\left|g_{T}-g_{0}\right|\rightarrow 0$ and using (d))
$\int\left|b_{T}g_{T}-b_{0}g_{0}\right|\leq\int\left|b_{T}g_{T}-b_{T}g_{0}\right|+\int\left|b_{T}g_{0}-b_{0}g_{0}\right|=\left|b_{T}\right|\int\left|g_{T}-g_{0}\right|+\left|b_{T}-b_{0}\right|\int\left|g_{0}\right|\longrightarrow 0.$
Letting $1/b_{T}=e^{\operatorname{\mathcal{PS}}_{T}\left(\theta_{T}\right)}T^{d/2}z_{T}$ and $1/b=\pi\left(\theta^{*}\right)\frac{(2\pi)^{d/2}}{\left|\operatorname{\Gamma}_{*}\right|^{1/2}}$, we have $b_{T}\rightarrow b$ by equation (18), and
$\int_{\mathbb{R}^{d}}\left|q_{T}(x)-\mathcal{N}\left(x\mid 0,\operatorname{\Gamma}_{*}^{-1}\right)\right|dx\underset{T\rightarrow\infty}{\longrightarrow}0,$
which proves equation (13) in our Theorem 2 and the concentration of $\pi_{T}$ around $\theta^{*}$ in equation (11) follows from the Lemma 28 in Miller (2021). To conclude the proof of the Theorem, next we show that the conditions mentioned above (a)–(d) hold.
Proof of (a)
Let us fix some $x\in\mathbb{R}^{d}$. For sufficiently large $T$, we have $|x|<\epsilon^{*}\sqrt{T}$, then,
$h_{T}(x)=\exp\left(-\frac{1}{2}x^{\prime}(\sum_{a\in\mathcal{A}}\frac{|T_{a}|}{T}A^{a}_{T})x\right)2\pi(\theta^{*})\longrightarrow\exp\left(-\frac{1}{2}x^{\prime}(\sum_{a\in\mathcal{A}}w_{a}A^{a}_{0})x\right)2\pi(\theta^{*})$
since $A^{a}_{T}\rightarrow A^{a}_{0}$ and $\frac{|T_{a}|}{T}\rightarrow w_{a}$ as $T\rightarrow\infty$ by assumptions A4 and A6 respectively and $|\mathcal{A}|<\infty$
Now, for $g_{T}$, first note that $\pi\left(\theta_{T}+x/\sqrt{T}\right)\rightarrow\pi\left(\theta_{0}\right)$ since $\pi$ is continuous at $\theta_{0}$ and $\theta_{T}\rightarrow\theta_{0},x/\sqrt{T}\rightarrow 0$. By assumption A4 for all action $a\in\mathcal{A}$, we have
$\operatorname{\mathcal{PS}}^{a}_{T}(\theta_{T}+x/\sqrt{T})-\operatorname{\mathcal{PS}}^{a}_{T}(\theta_{T})=\frac{1}{2}\frac{|T_{a}|}{T}x^{\prime}\operatorname{\Gamma}^{a}_{T}x+|T_{a}|r^{a}_{T}(x/\sqrt{T})\longrightarrow\frac{1}{2}w_{a}x^{\prime}\operatorname{\Gamma}^{a}_{*}x,$ (19)
since $\operatorname{\Gamma}^{a}_{T}\rightarrow\operatorname{\Gamma}^{a}_{*}$ and for all $T$ sufficiently large (to ensure that $|x/\sqrt{T}|<\varepsilon_{0}$ and the assumed bound on $r^{a}_{T}$ holds),
$\left||T_{a}|r^{a}_{T}(x/\sqrt{T})\right|\leq|T_{a}|c_{0}|x/\sqrt{T}|^{3}=\frac{|T_{a}|}{T}\frac{c_{0}|x|^{3}}{\sqrt{T}}\rightarrow 0$
as $T\rightarrow\infty$. Next, let us consider
$g_{T}(x)$ $=\exp[-\sum_{a}(\operatorname{\mathcal{PS}}^{a}_{T}(\theta_{T}+x/\sqrt{T})-\operatorname{\mathcal{PS}}^{a}_{T}(\theta_{T}))]\pi(\theta_{T}+x/\sqrt{T})$
$=\prod_{a}[\exp[-(\operatorname{\mathcal{PS}}^{a}_{T}(\theta_{T}+x/\sqrt{T})-\operatorname{\mathcal{PS}}^{a}_{T}(\theta_{T}))](\pi(\theta_{T}+x/\sqrt{T}))^{\frac{1}{|\mathcal{A}|}}]$
@@
$=\exp\left(-\frac{1}{2}x^{\mathrm{T}}\operatorname{\Gamma}_{*}x\right)\pi\left(\theta^{*}\right)=g_{0}(x).$
Hence, $g_{T}(x)\rightarrow g_{0}(x)$ as $T\rightarrow\infty$, which completes the proof of (a).
Proof of (b)
By the definition of $h_{T}$ and $A_{T}=\sum_{a}\frac{|T_{a}|}{T}A_{T}^{a}$, letting $B_{T}=B_{\varepsilon^{*}\sqrt{T}}(0)$,
@@
Since $A^{a}_{T}\rightarrow A^{a}_{0}$ as $T\rightarrow\infty$ and $A^{a}_{0}$ is positive definite, then for all $T$ sufficiently large, $A^{a}_{T}$ is also positive definite for all $a\in\mathcal{A}$. Then the first term equals
$2\pi\left(\theta_{0}\right)\frac{(2\pi)^{d/2}}{\left|\sum_{a}\frac{|T_{a}|}{T}A^{a}_{T}\right|^{1/2}}{P}\left(\left|\left(\sum_{a}\frac{|T_{a}|}{T}A^{a}_{T}\right)^{-1/2}Z\right|<\varepsilon^{*}\sqrt{T}\right)\longrightarrow 2\pi\left(\theta_{0}\right)\frac{(2\pi)^{d/2}}{\left|\sum_{a}w_{a}A^{a}_{0}\right|^{1/2}}=\int h_{0},$
where $Z\sim\mathcal{N}(0,I)$.
The second term goes to zero, since it is nonnegative and upper bounded by
[…]
using the fact that $\pi\left(\theta_{T}+x/\sqrt{T}\right)T^{-d/2}$ is the density of $X=\sqrt{T}\left(\theta-\theta_{T}\right)$ when $\theta\sim\pi$.
Proof of (c)
We consider $T$ sufficiently large such that $\left|\theta_{T}-\theta^{*}\right|<\varepsilon^{*}$ (using Lemma  1), the bound on $r^{a}_{T}$ applies, and $\inf_{\theta\in B_{\varepsilon^{*}}\left(\theta_{T}\right)^{c}}\frac{1}{|T_{a}|}(\operatorname{\mathcal{PS}}^{a}_{T}(\theta)-\operatorname{\mathcal{PS}}^{a}_{T}\left(\theta_{T}\right))>\delta_{a}/2$ for all $a\in\mathcal{A}$ by assumption A5.
If $|x|\geq\varepsilon^{*}\sqrt{T}$, then $\left(\operatorname{\mathcal{PS}}^{a}_{T}\left(\theta_{T}+x/\sqrt{T}\right)-\operatorname{\mathcal{PS}}^{a}_{T}\left(\theta_{T}\right)\right)>|T_{a}|\delta_{a}/2$ for all $a\in\mathcal{A}$, and thus,
$\operatorname{\mathcal{PS}}_{T}\left(\theta_{T}+x/\sqrt{T}\right)-\operatorname{\mathcal{PS}}_{T}\left(\theta_{T}\right)>\frac{T}{2}\sum_{a}\frac{|T_{a}|}{T}\delta_{a}=\frac{T\delta}{2},$
which implies,
$g_{T}(x)\leq e^{-T\delta/2}\pi\left(\theta_{T}+x/\sqrt{T}\right)=h_{T}(x).$
On the other hand, if $|x|<\varepsilon^{*}\sqrt{T}$, then $\pi\left(\theta_{T}+x/\sqrt{T}\right)\leq 2\pi\left(\theta_{0}\right)$ (by our choice of $\varepsilon^{*}$, since $|(\theta_{T}+x/\sqrt{T})-\theta_{0}|\leq|\theta_{T}-\theta_{0}|+|x/\sqrt{T}|<2\varepsilon^{*})$. Further, $\left||T_{a}|r^{a}_{T}(x/\sqrt{T})\right|\leq\frac{|T_{a}|}{T}\frac{c_{0}|x|^{3}}{\sqrt{T}}\leq\frac{|T_{a}|}{T}c_{0}\varepsilon^{*}|x|^{2}\leq\frac{1}{2}\frac{|T_{a}|}{T}\alpha^{a}|x|^{2}$, by the fact that $|x/\sqrt{T}|<\varepsilon^{*}<\varepsilon_{0}$ and $\varepsilon^{*}<\alpha^{a}/\left(2c_{0}\right)$ for all $a\in\mathcal{A}$. Using the assumption A4 we get,
$\operatorname{\mathcal{PS}}_{T}\left(\theta_{T}+x/\sqrt{T}\right)-\operatorname{\mathcal{PS}}_{T}\left(\theta_{T}\right)$ $=\sum_{a}\left[\frac{1}{2}\frac{|T_{a}|}{T}x^{\prime}\operatorname{\Gamma}^{a}_{T}x+|T_{a}|r^{a}_{T}(x/\sqrt{T})\right]$
$\geq\sum_{a}\left[\frac{1}{2}\frac{|T_{a}|}{T}x^{\prime}\operatorname{\Gamma}^{a}_{T}x-\frac{1}{2}\frac{|T_{a}|}{T}\alpha^{a}x^{\prime}x\right]$
$=\frac{1}{2}x^{\prime}\left(\sum_{a}\frac{|T_{a}|}{T}(\operatorname{\Gamma}^{a}_{T}-\alpha^{a}I_{d})\right)x=\frac{1}{2}x^{\prime}\left(\sum_{a}\frac{|T_{a}|}{T}A^{a}_{T}\right)x=\frac{1}{2}x^{\prime}A_{T}x,$
implying,
$g_{T}(x)\leq\exp\left(-\frac{1}{2}x^{\prime}A_{T}x\right)2\pi\left(\theta_{0}\right)=h_{T}(x).$
Proof of (d)
Since $\operatorname{\Gamma}^{a}_{*}$ and $A^{a}_{0}$ are positive definite for all $a\in\mathcal{A}$, $\int g_{0}$ and $\int h_{0}$ are finite. By (b) and (c), since $\int h_{T}\rightarrow\int h_{0}<\infty$, we have $\int g_{T}\leq\int h_{T}<\infty$ for all $T$ sufficiently large. Measurability of $g_{T}$ and $h_{T}$ follows from measurability of $\operatorname{\mathcal{PS}}_{T}$ and $\pi$. This concludes our proof.
A.2 Approximation error of the Q function

### v2

@@ A.1 Proof of Lemma 1
$\lim_{T\to\infty}\frac{1}{T}\widetilde{\operatorname{\mathcal{PS}}}_{T}(\theta)=\lim_{T\to\infty}\frac{1}{T}\widetilde{\operatorname{\mathcal{PS}}}_{T}(\mathbb{M}_{\theta},\mathbb{M}_{0})$ $=\lim_{T\to\infty}\frac{1}{T}\left(D_{T}(\theta)+\widetilde{\operatorname{\mathcal{PS}}}_{T}(\mathbb{M}_{0},\mathbb{M}_{0})\right)$
$=D^{*}(\theta)+c^{*}=\operatorname{\mathcal{PS}}^{*}(\theta).$
A.2 Proof of Lemma 2
Lemma 2 (Uniform law of large numbers) Under assumptions A2 and A3, a uniform law of large numbers (ULLN) holds, which implies with probability $1$ under $\mathbb{M}_{0}$,
$\sup_{\theta\in\Theta}|\operatorname{\mathcal{PS}}_{T}(\theta)-\widetilde{\operatorname{\mathcal{PS}}}_{T}(\theta)|\to 0\text{ as }T\to\infty.$
Proof  We obtain a ULLN from the stationarity, mixing and moment boundedness conditions in assumptions A2 and A3 from a result in Pötscher and Prucha (1989) as adapted by Pacchiardi et al. (2024a). According to the ULLN, for all actions $a\in\mathcal{A}$ with probability $1$ under $\mathbb{M}_{0}$,
$\sup_{\theta\in\Theta}|\operatorname{\mathcal{PS}}_{T}^{a}(\theta)-\widetilde{\operatorname{\mathcal{PS}}}_{T}^{a}(\theta)|\to 0,$
where $\widetilde{\operatorname{\mathcal{PS}}}_{T}^{a}(\theta)=\mathbb{E}_{H_{T}^{a}\sim\mathbb{M}_{0}}\operatorname{\mathcal{PS}}(\mathbb{M}_{\theta},H_{T}^{a})$ denotes the expected prequential score calculated with respect to the observations where action $a$ was chosen. Then, using the triangle inequality we get with probability $1$ under $\mathbb{M}_{0}$,
$\sup_{\theta\in\Theta}|\operatorname{\mathcal{PS}}_{T}(\theta)-\widetilde{\operatorname{\mathcal{PS}}}_{T}(\theta)|$ $=\sup_{\theta\in\Theta}|\sum_{a\in\mathcal{A}}\operatorname{\mathcal{PS}}_{T}^{a}(\theta)-\sum_{a\in\mathcal{A}}\widetilde{\operatorname{\mathcal{PS}}}_{T}^{a}(\theta)|$
$\leq\sup_{\theta\in\Theta}\sum_{a\in\mathcal{A}}|\operatorname{\mathcal{PS}}_{T}^{a}(\theta)-\widetilde{\operatorname{\mathcal{PS}}}_{T}^{a}(\theta)|\to 0.$ (17)
A.3 Proof of Corollary 3
Corollary 3 Under assumptions A1-A3, it follows from Lemma 1 and Lemma 2 that,
$\sup_{\theta\in\Theta}|\frac{1}{T}\operatorname{\mathcal{PS}}_{T}(\theta)-{\operatorname{\mathcal{PS}}}^{*}(\theta)|\to 0\text{ as }T\to\infty,$
with probability $1$ under $\mathbb{M}_{0}$.
Proof  Under assumption A1, from Lemma 1, for a fixed $\epsilon>0$ there exists a $T_{1}(\epsilon)$ such that for all $T>T_{1}(\epsilon)$ with probability $1$ under $\mathbb{M}_{0}$,
$|\frac{1}{T}\widetilde{\operatorname{\mathcal{PS}}}_{T}(\theta)-\operatorname{\mathcal{PS}}^{*}(\theta)|<\epsilon/2,\text{ for all }\theta\in\Theta$
$\implies\sup_{\theta\in\Theta}|\frac{1}{T}\widetilde{\operatorname{\mathcal{PS}}}_{T}(\theta)-\operatorname{\mathcal{PS}}^{*}(\theta)|<\epsilon/2.$ (18)
Similarly, under assumptions A2 and A3, Lemma 2 implies that for the fixed $\epsilon>0$ there exists a $T_{2}(\epsilon)$ such that for all $T>T_{2}(\epsilon)$ with probability $1$ under $\mathbb{M}_{0}$,
$\sup_{\theta\in\Theta}|{\operatorname{\mathcal{PS}}}_{T}(\theta)-\widetilde{\operatorname{\mathcal{PS}}}(\theta)|<\epsilon T/2.$ (19)
On combining equations A.3 and 19, for any $T>\max\{T_{1}(\epsilon),T_{2}(\epsilon)\}$
$\sup_{\theta\in\Theta}|\frac{1}{T}{\operatorname{\mathcal{PS}}}_{T}(\theta)-\operatorname{\mathcal{PS}}^{*}(\theta)|$ $\leq\sup_{\theta\in\Theta}\left\{|\frac{1}{T}{\operatorname{\mathcal{PS}}}_{T}(\theta)-\frac{1}{T}\widetilde{\operatorname{\mathcal{PS}}}_{T}(\theta)|+|\frac{1}{T}\widetilde{\operatorname{\mathcal{PS}}}_{T}(\theta)-\operatorname{\mathcal{PS}}^{*}(\theta)|\right\}$
$<\epsilon/2+\epsilon/2=\epsilon,$ (20)
with probability $1$ under $\mathbb{M}_{0}$. Since equation A.3 holds for an arbitrary $\epsilon>0$, it follows that, with probability $1$ under $\mathbb{M}_{0}$,
$\sup_{\theta\in\Theta}|\frac{1}{T}{\operatorname{\mathcal{PS}}}_{T}(\theta)-\operatorname{\mathcal{PS}}^{*}(\theta)|\to 0\text{ as }T\to\infty.$
A.4 Proof of Lemma 4
Lemma 4 (Asymptotic consistency) Under assumptions A1-A5, as $T\to\infty$, we have with probability $1$ under $\mathbb{M}_{0}$,
$d(\hat{\theta}_{T},\theta^{*})\to 0.$
We present here a proof of the lemma based on Theorem 5.1 in Skouras (1998) as adapted by Pacchiardi et al. (2024a).
Proof  From assumption A5, for a fixed $\epsilon>0$ it is possible to find a $\delta(\epsilon)>0$ such that,
$\min_{\theta:d(\theta,\theta^{*})\geq\epsilon}{\operatorname{\mathcal{PS}}}^{*}(\theta)-{\operatorname{\mathcal{PS}}}^{*}(\theta^{*})=\delta(\epsilon),$ (21)
with probability $1$ under $\mathbb{M}_{0}$.
Due to Corollary 3, with probability $1$ under $\mathbb{M}_{0}$, there exists a $T_{1}(\delta(\epsilon))$ such that for all $T>T_{1}(\delta(\epsilon))$
$|\frac{1}{T}{\operatorname{\mathcal{PS}}}_{T}(\theta^{*})-\operatorname{\mathcal{PS}}^{*}(\theta^{*})|<\delta(\epsilon)/2,$
which implies
$\operatorname{\mathcal{PS}}^{*}(\theta^{*})$ $>\frac{1}{T}{\operatorname{\mathcal{PS}}}_{T}(\theta^{*})-\delta(\epsilon)/2$
$\geq\frac{1}{T}{\operatorname{\mathcal{PS}}}_{T}(\hat{\theta}_{T})-\delta(\epsilon)/2,$ (22)
where the second inequality is from the definition of $\hat{\theta}_{T}$.
On exploiting Corollary 3 once again, we can define a $T_{2}(\delta(\epsilon))$ such that for all $T>T_{2}(\delta(\epsilon))$
$|\frac{1}{T}{\operatorname{\mathcal{PS}}}_{T}(\hat{\theta}_{T})-\operatorname{\mathcal{PS}}^{*}(\hat{\theta}_{T})|<\delta(\epsilon)/2,$ (23)
with probability $1$ under $\mathbb{M}_{0}$.
Then, with probability $1$ under $\mathbb{M}_{0}$, for all $T>\max\{T_{1}(\delta(\epsilon)),T_{2}(\delta(\epsilon))\}$
$\operatorname{\mathcal{PS}}^{*}(\hat{\theta}_{T})-\operatorname{\mathcal{PS}}^{*}(\theta^{*})$ $=\operatorname{\mathcal{PS}}^{*}(\hat{\theta}_{T})-\frac{1}{T}{\operatorname{\mathcal{PS}}}_{T}(\hat{\theta}_{T})+\frac{1}{T}{\operatorname{\mathcal{PS}}}_{T}(\hat{\theta}_{T})-\operatorname{\mathcal{PS}}^{*}(\theta^{*})$
$<\delta(\epsilon)/2+\delta(\epsilon)/2=\delta(\epsilon)$ (24)
from equation A.4 and equation 23.
Note that equation A.4 ensures that the difference considered in equation 21 is smaller than $\delta(\epsilon)$ when $\theta=\hat{\theta}_{T}$. However, by equation 21, the same difference is at least $\delta(\epsilon)$ for all $\theta$ that are outside the $\epsilon$-radius ball around $\theta^{*}$. This implies that, $\hat{\theta}_{T}$ must lie inside the $\epsilon$-radius ball, meaning that $d(\hat{\theta}_{T},\theta^{*})<\epsilon$ with probability $1$ under $\mathbb{M}_{0}$. Since this is true for any $\epsilon>0$, it follows that, with probability $1$ under $\mathbb{M}_{0}$
$d(\hat{\theta}_{T},\theta^{*})\to 0\text{ as }T\to\infty.$
A.5 Sketch proof of BvM theorem
Here we present a sketch proof for the BvM theorem (Theorem 5) based on some results from Miller (2021).
Proof  From Lemma 4 we have with probability one under $\mathbb{M}_{0}$, as $T\to\infty$,
$d(\hat{\theta}_{T},\theta^{*})\to 0,$
which proves the existence of a sequence of estimators $\{\hat{\theta}_{T}\}_{T=1}^{\infty}$ such that as $T\to\infty$, with probability one under $\mathbb{M}_{0}$, $d(\hat{\theta}_{T},\theta^{*})\to 0$ under the assumptions A1-A5.
By Theorem 6 from Miller (2021), using the above sequence of estimators $\{\hat{\theta}_{T}\}_{T=1}^{\infty}$ along with the assumptions A7-A9, $\operatorname{\mathcal{PS}}_{T}(\theta)$ can be represented as,
$\operatorname{\mathcal{PS}}_{T}(\theta)=\operatorname{\mathcal{PS}}_{T}(\hat{\theta}_{T})+\frac{1}{2}(\theta-\hat{\theta}_{T})^{\prime}H_{T}(\theta-\hat{\theta}_{T})+r_{T}(\theta-\hat{\theta}_{T})$ (25)
where $H_{T}=\operatorname{\mathcal{PS}}_{T}^{\prime\prime}(\hat{\theta}_{T})\in\mathbb{R}^{p\times p}$ is symmetric and $H_{T}\to H^{*}$. There exists $\epsilon_{0},c_{0}>0$ such that, for all $T$ sufficiently large, for all $\theta\in B_{\epsilon_{0}}(\mathbf{0})$, we have $|r_{T}(\theta)|\leq c_{0}|\theta|^{3}.$
Finally, on using the Taylor series expansion of the function $\operatorname{\mathcal{PS}}_{T}(\theta)$ in equation 25, together with the assumptions A6 and A10, the Theorem 5 holds (by Theorem 4 from Miller (2021)).
For completeness, we report the full statements of Theorem 4 and Theorem 6 from Miller (2021) below as Theorem 7 and Theorem 8, respectively.
Theorem 7 (Theorem 4 from Miller (2021))
Fix $\theta^{*}\in\mathbb{R}^{p}$ and let $\pi:\mathbb{R}^{p}\rightarrow\mathbb{R}$ be a probability density with respect to Lebesgue measure such that $\pi$ is continuous at $\theta^{*}$ and $\pi\left(\theta^{*}\right)>0$. Let $\operatorname{\mathcal{PS}}_{T}:\mathbb{R}^{p}\rightarrow\mathbb{R}$ for $T\in\mathbb{N}$ and assume:
  1. M1
    $L_{T}$ can be represented as
    $\operatorname{\mathcal{PS}}_{T}(\theta)=\operatorname{\mathcal{PS}}_{T}(\hat{\theta}_{T})+\frac{1}{2}(\theta-\hat{\theta}_{T})^{\prime}H_{T}(\theta-\hat{\theta}_{T})+r_{T}(\theta-\hat{\theta}_{T})$
    where $\hat{\theta}_{T}\in\mathbb{R}^{p}$ such that $\hat{\theta}_{T}\rightarrow\theta^{*},H_{T}\in\mathbb{R}^{p\times p}$ symmetric such that $H_{T}\rightarrow H^{*}$ for some positive definite $H^{*}$, and $r_{T}:\mathbb{R}^{p}\rightarrow\mathbb{R}$ has the following property: there exist $\varepsilon_{0},c_{0}>0$ such that for all $T$ sufficiently large, for all $\theta\in B_{\varepsilon_{0}}(\mathbf{0})$, we have $\left|r_{T}(\theta)\right|\leq c_{0}|\theta|^{3};$
  2. M2
    For any $\varepsilon>0,\liminf_{T}\inf_{\theta\in B_{\varepsilon}\left(\hat{\theta}_{T}\right)^{c}}\left(\operatorname{\mathcal{PS}}_{T}(\theta)-\operatorname{\mathcal{PS}}_{T}\left(\hat{\theta}_{T}\right)\right)>0$,
@@ A.5 Sketch proof of BvM theorem
which means, $\pi_{GP}(\theta|y^{T})$ concentrates at $\theta^{*}$;
$z_{T}\approx\frac{\exp(-A_{T}L_{T}(\hat{\theta}_{T}))\pi(\theta^{*})}{\mid\det H^{*}\mid^{1/2}}\left(\frac{2\pi}{T}\right)^{d/2}$
as $T\to\infty$ (Laplace approximation), and letting $q_{T}$ be the density of $\sqrt{T}(\theta-\hat{\theta}_{T})$ when $\theta\sim\pi_{GP}(\theta|y^{T})$,
$\int_{\Theta}\left|q_{T}(\theta)-\mathcal{N}(\theta\mid\mathbf{0},{H^{*}}^{-1})\right|d\theta\underset{T\rightarrow\infty}{\longrightarrow}0,$
which implies, $q_{T}$ converges to $\mathcal{N}(\mathbf{0},{H^{*}}^{-1})$ in total variation.
Theorem 8 (Theorem 6 from Miller (2021))
Let $E\subseteq\mathbb{R}^{p}$ be open and convex, and let $\theta^{*}\in E$. Let $L_{T}:E\rightarrow\mathbb{R}$ have continuous third derivatives, and assume:
  1. M3
    there exist $\hat{\theta}_{T}\in E$ such that $\hat{\theta}_{T}\rightarrow\theta^{*}$ and $L_{T}^{\prime}(\hat{\theta}_{T})=0$ for all $T$ sufficiently large,
  2. M4
    $L_{T}^{\prime\prime}\left(\theta^{*}\right)\rightarrow H^{*}$ as $T\rightarrow\infty$ for some positive definite $H^{*}$, and
  3. M5
    $L_{T}^{\prime\prime\prime}$ is uniformly bounded;
then, letting $H_{T}=L_{T}^{\prime\prime}(\hat{\theta}_{T})$, condition M1 is satisfied for all $T$ sufficiently large.
Appendix B Approximation error of the Q function